OpenStudy (icalibear):

Systems of equations help please A student took a science test and a math test. They scored 10 points more on the math test than they did on the science test. The average score of the two tests was 75. The average score is given by x + y / 2. x = science, y = math. what did they score on the math test

jimthompson5910 (jim_thompson5910):

x = science, y = math "They scored 10 points more on the math test than they did on the science test" so how can you use that sentence to tie x and y together? what equation can we form?

OpenStudy (icalibear):

y = x + 10 and x = y - 10?

jimthompson5910 (jim_thompson5910):

they're the same equation, so let's go with y = x + 10

OpenStudy (icalibear):

okay

jimthompson5910 (jim_thompson5910):

Then it says "The average score of the two tests was 75" we will use the average formula A = (x+y)/2 to get \[\Large A = \frac{x+y}{2}\] \[\Large 75 = \frac{x+y}{2}\] \[\Large 75 = \frac{x+x+10}{2}\] In line 3, I replaced y with x+10 since y = x+10 Now solve for x

OpenStudy (icalibear):

i got 15

jimthompson5910 (jim_thompson5910):

that's too small

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (icalibear):